Pyramid

Given an integer N, write a program to print a pyramid of N rows as shown below.

Input

The first line of input is an integer N.

Explanation

In the example, the number of rows in the pyramid is

5.So, the output should be


. . . . 0 . . . .

. . . 0 0 0 . . .

. . 0 0 0 0 0 . .

. 0 0 0 0 0 0 0 .

0 0 0 0 0 0 0 0 0


Sample Input 1

5

Sample Output 1

. . . . 0 . . . .

. . . 0 0 0 . . .

. . 0 0 0 0 0 . .

. 0 0 0 0 0 0 0 .

0 0 0 0 0 0 0 0 0

Sample Input 2

6

Sample Output 2

. . . . . 0 . . . . .

. . . . 0 0 0 . . . .

. . . 0 0 0 0 0 . . .

. . 0 0 0 0 0 0 0 . .

. 0 0 0 0 0 0 0 0 0 .

0 0 0 0 0 0 0 0 0 0 0




1
Expert's answer
2021-09-06T23:58:16-0400
N = int(input("Enter an interger: "))
for i in range(N):
    print('. ' * (N - i - 1), end='')
    print('0 ' * (2 * i + 1), end='')
    print('. ' * (N - i - 1), end='')
    print()